OpenStudy (anonymous):

THE PERIOD OF THIS COS FUNCTION? H(t) = 20 cos(pi/15)t + 30

OpenStudy (jdoe0001):

\(\bf H(t)=20cos\left(\frac{\pi}{15}t\right)+30\quad ?\)